# StepProof: Step-by-step verification of natural language mathematical proofs

### v1

@@ 2 Related Work
Autoformalization: The definition of automatic formalization is very broad, but can be roughly seen as understanding and extracting translation from natural language to obtain the required structured data or formal language, such as entity relation extraction (nguyen2015relation, ). Early automatic formalization work involved extracting logical propositions from natural language in addition to entity relation extraction (singh2020exploring, ; lu2022parsing, ). However, the main problem in this kind of work is that the extracted logical propositions lack corresponding symbols for derivation and application, so the output of the automatic formalization output cannot be directly applied. With the launch of pre-trained language models such as transformer and BERT, language models have a stronger understanding ability. wang2020exploration conducted an early automatic formalization attempt for the theorem prover Mizar. However, due to the limited size and training expectations of the models at that time, the effects they could achieve were very limited. With the rapid expansion of the scale of models and predictions, many new and better performance automatic formalization work has emerged, such as the Majority Voting method by lewkowycz2022solvingquantitativereasoningproblems , the DSP method by jiang2022draft ., and the method proposed by zhou2024don , in DTV (Don’t Trust: Verify). They further improved the automatic formalization of the system by combining large models with some syntax-modifying filters. However, on the one hand, all these works are tested in the closed-source large model Minerva, which lacks the testing work of open source model and small model. Meanwhile, all these works adopt the strategy of FULL-PROOF, which has poor controllability for the formal output of the model, and it is difficult to locate the error point of non-formal proof. Qinghua et al. proposed a problem location method based on the FULL-PROOF strategy in SlideRule. However, this method relies heavily on the format and quality of the generated formal content, so it cannot achieve 100% problem locations detected. To solve these problems, we propose StepProof, an automatic formalization strategy that can realize sentence-level verification, to realize the verification of natural language mathematical proofs. Moreover, LEGO-Prover (wang2023lego, ) proposed a new methodology to decompose the whole proof into several sub-proofs. Although the idea of step-proof is taking shape, it still requires some extra generation of the sub-proof formal statement generation, which increases the error probability of formalization.
@@ 4.1 Experiment Setup
In the baseline tests, we compared the multi-attempt test results of GSM8K by lewkowycz2022solvingquantitativereasoningproblems in Majority Voting and zhou2024don in DTV with our results. We evaluated the performance based on the number of attempts and multi-round proof pass rate, allowing up to 10 retries for each failed step in each formalization attempt.

### v2

@@ 2 Related Work
Autoformalization: The definition of automatic formalization is very broad, but can be roughly seen as understanding and extracting translation from natural language to obtain the required structured data or formal language, such as entity relation extraction (nguyen2015relation, ). Early automatic formalization work involved extracting logical propositions from natural language in addition to entity relation extraction (singh2020exploring, ; lu2022parsing, ). However, the main problem in this kind of work is that the extracted logical propositions lack corresponding symbols for derivation and application, so the output of the automatic formalization output cannot be directly applied. With the launch of pre-trained language models such as transformer and BERT, language models have a stronger understanding ability. Wang et al.(wang2020exploration, ) conducted an early automatic formalization attempt for the theorem prover Mizar. However, due to the limited size and training expectations of the models at that time, the effects they could achieve were very limited. With the rapid expansion of the scale of models and predictions, many new and better performance automatic formalization work has emerged, such as the Majority Voting method by Lewkowycz et al.(lewkowycz2022solvingquantitativereasoningproblems, ), the DSP method by Jiang et al.(jiang2022draft, )., and the method proposed by Zhou et al.(zhou2024don, ), in DTV (Don’t Trust: Verify). They further improved the automatic formalization of the system by combining large models with some syntax-modifying filters. However, on the one hand, all these works are tested in the closed-source large model Minerva, which lacks the testing work of open source model and small model. Meanwhile, all these works adopt the strategy of FULL-PROOF, which has poor controllability for the formal output of the model, and it is difficult to locate the error point of non-formal proof. Qinghua et al. proposed a problem location method based on the FULL-PROOF strategy in SlideRule. However, this method relies heavily on the format and quality of the generated formal content, so it cannot achieve 100% problem locations detected. To solve these problems, we propose StepProof, an automatic formalization strategy that can realize sentence-level verification, to realize the verification of natural language mathematical proofs. Moreover, LEGO-Prover (wang2023lego, ) proposed a new methodology to decompose the whole proof into several sub-proofs. Although the idea of step-proof is taking shape, it still requires some extra generation of the sub-proof formal statement generation, which increases the error probability of formalization.
@@ 4.1 Experiment Setup
In the baseline tests, we compared the multi-attempt test results of GSM8K by Lewkowycz et al.(lewkowycz2022solvingquantitativereasoningproblems, ) in Majority Voting and Zhou et al.(zhou2024don, ) in DTV with our results. We evaluated the performance based on the number of attempts and multi-round proof pass rate, allowing up to 10 retries for each failed step in each formalization attempt.